OpenStudy (anonymous):

Solve for the variable : -m/4.2 = 5

OpenStudy (anonymous):

Multiply 5 by both sides?

OpenStudy (shane_b):

Multiply both sides by 4.2 so that you can get rid of the fraction

OpenStudy (shane_b):

\[\frac{-m}{4.2}=5\]\[(4.2)\frac{-m}{4.2}=5(4.2)\]

OpenStudy (shane_b):

That will leave you with -m=21. So m will equal -21